OpenStudy (anonymous):

I need help expanding this binomial: (3x+6)^3

OpenStudy (anonymous):

(3x+6)^3 = 3^3x^3+6^3 = 27x^3+216

OpenStudy (nincompoop):

familiar with pascal's triangle?

OpenStudy (anonymous):

indices

OpenStudy (nincompoop):

http://www.purplemath.com/modules/binomial.htm

OpenStudy (anonymous):

that's different

OpenStudy (nincompoop):

uh no...

OpenStudy (nincompoop):

\[ (3x+6)^3 = (3x+6)(3x+6)(3x+6) \]

OpenStudy (anonymous):

@chamyiyao was correct on this one.

OpenStudy (nincompoop):

laughing out loud okay then

OpenStudy (nincompoop):

good luck with your studies

OpenStudy (anonymous):

thanks @nincompoop Not sure why but the last term was what I needed for the question, and the correct answer was 216. I admire your sassy attitude though ;)